OpenStudy (anonymous):

Aaron is painting all 4 walls of his room. Each wall is 8 feet by 12 feet. One wall has a 4 foot by 5 foot window and one wall has a 3 foot by 7.5 foot door. Neither of these will be painted. Aaron only wants to paint one coat on each wall. He learns that each gallon of paint will cover 125 ft2. He can only buy entire gallons of paint. How many gallons of paint will he need?

OpenStudy (anonymous):

4 walls (8ft x 12 ft) (one coat of paint on each wall) one wall has a 4ft x 5ft window (not painted) one wall has a 3 ft x 7.5 ft door(not painted) 1 gallon of paint = 125ft^2 first calculte 4 walls = 8 x 12 = 96ft2 x 4 walls = 384ft2

OpenStudy (anonymous):

next deduct space for 4ft x 5ft window and 3ft x 7.5 ft door = 20ft2 + 22.5ft2 = 42.5ft2 deducted from total room = 384ft2 - 42.5ft2 = 341.5ft2

OpenStudy (anonymous):

each gallon of paint will cover 125ft2 so therefore, dividing this into the amount of space in our room, gives: 341.5ft2/125ft2 = 2.732 cans. ~Since a gallon will cover 125ft2 and he can only buy gallons,, he therefore must buy 3 x 125ft gallons to make sure he covers the whole area of the room Answer: 3 gallons of paint